How do I calculate the zeros of the following function?
F(x) = x^3+x^2-49x-49,
Given (x-1) is a factor of the polynomial.

Answer by ankor@dixie-net.com(22740)

x^3+x^2-49x-49,
group factor
(x^3 + x^2) - (49x - 49) = 0
Factor out the common term
x^2(x + 1) - 49(x + 1) = 0
Factor out (x+1)
(x + 1)(x^2 - 49) = 0
Factor (x^2 - 49) as the difference of squares
(x -7)(x + 7) = 0
Three zeros
x = 1
x = 7
x = -7